# Hierarchies of direct powers, ultrapowers and cumulative powers

### 16 Mar 2026 v1

@@ 2.1. A concrete example
For $f\in\mathcal{F}^{*}_{1}(\mathbb{N})=\mathbb{N}^{\mathbb{N}}$, let $f^{+}$ be defined as $\forall j\in\mathbb{N}\big(g^{+}(j)=\mathrm{s}(g(j))\big)$. Consider the function $\mathrm{s}_{1}$, the extension of the successor function $\mathrm{s}$ to $\mathcal{F}_{1}(\mathbb{N})$:
@@
$\forall j\in\mathcal{F}_{1}(\mathbb{N})\big(a(j)=f(j)+_{1}3\big)$,
@@
$a(j)=\begin{cases}j^{+}+_{1}3,\ \text{if}\ j\in\mathbb{N}^{\mathbb{N}}\\ \mathrm{s}(1)+3,\ \text{if}\ j\in\mathbb{N}\end{cases}$
@@
In order to have a visual representation of the this sum we may take an order of $\mathcal{F}_{1}(\mathbb{N})$ such that $\mathbb{N}$ is the initial segment, and hence we may represent an element of $\mathcal{F}_{2}(\mathbb{N})$ by a column vector with the $j$-th coordinate being its value at $j$. Then the first vector below represents $\mathrm{s}_{1}$, and the sum $\mathrm{s}_{1}+_{2}3$ can be depicted as follows.
@@ 2.2. Preservation and properties of cumulative powers
In [13], Horn showed that any Horn sentence is preserved by direct products. In [4], Chang and Morel showed Horn sentences are also preserved by reduced products produced by the Frechet filter, and as expressed in [8], Chang further generalized the result, showing that, in fact, any proper (that is, defined by proper filters) reduced product preserves Horn sentences.
A class of structures defined by a set of sentences is a elementary class, and if all those sentences are Horn sentences, call it a Horn class. The results above mentioned therefore say that any Horn class is closed by proper reduced (and thus direct) products. Chang and Morel showed in [4] that not every elementary class closed under direct products is a Horn class, but conjectured that any elementary class closed under reduced products is indeed a Horn class. In [14], Keisler proved the conjecture under the assumption the continuum hypothesis holds for at least one cardinal, whereas in [10], Galvin strengthened the result by eliminating that assumption. Briefly:
@@
Let the occurrence of an atom in a formula be positive if in the prenex conjunctive normal form of that formula the atom is preceded by an even number of negations, and let it be negative otherwise. As such, one might suspect the fragment preserved by cumulative powers is the class of Horn sentences without positive equalities, for, as we have seen, there is an imbalance in how the interpretation of constants, functions and other relations are defined in cumulative powers (that is, hereditarily), and how equality is defined (that is, as true identity, rather than hereditary identity). In that case, cumulative powers would preserve strictly less first-order sentences than reduced products. However, consider the sentence $\forall x\forall y\exists z(x+y=z)$. As one may check on our previous example of a hierarchy given by the natural numbers, that sentence is preserved and has a positive occurrence of equality. Therefore, the matter is not as simple. Furthermore, as we may comparatively see on the definitions of cumulative powers and direct powers, the former are much closer to latter than to reduced products or reduced powers. Truth of formulas for functional elements of cumulative powers accord to how their truth are defined in direct powers. Therefore, if we want to characterize the fragment of first-order theory preserved by cumulative powers, our route should be to look at what fragment is preserved by direct powers.
In [19], Weinstein showed it is possible to give a syntactic – although not simple – characterization of that fragment. To expose it, some definitions are needed.
@@
Let $\Phi$ be a finite autonomous set of formulas closed under $\bigwedge$ and $\bigvee$ up to equivalence. Let $x$ be any variable and let $\varphi$, $\varphi^{\prime}$, and $\psi$ be formulas in $\bigvee\bigwedge\exists xF$. Fix $Q\in\{\exists,\forall\}$. Suppose for every disjunct $\varphi^{\prime}$ of $\varphi$ and every disjunct $\psi^{\prime}$ of $\psi$, there is a disjunct $\gamma^{\prime}$ of $\gamma$ such that for every conjunct $Qx\gamma^{\prime\prime}$ of $\gamma^{\prime}$, there are conjuncts $Qx\varphi^{\prime\prime}$ of $\varphi^{\prime}$ and $qx\psi^{\prime\prime}$ of $\psi^{\prime}$ such that $\varphi^{\prime\prime}\times\psi^{\prime\prime}\Rightarrow\gamma^{\prime\prime}$. Then, $\varphi\times\psi\Rightarrow\gamma$. Moreover, if $\varphi$ and $\psi$ are elements of $\bigvee\bigwedge^{*}Qx\Phi$, then the supposition is necessary and sufficient for $\varphi\times\psi\Rightarrow\gamma$.
@@
Notice $R$ is a computable predicate, as, given formulas $\varphi$, $\psi$ and $\gamma$, we can take $\Phi$ to be the set of atomic formulas occurring in $\varphi$, $\psi$ and $\gamma$, and it suffices to check for $n$ that is as large as the greatest between the quantifier ranks of those formulas.
@@
Notice we may replace products with powers in the above corollary, that is:
@@
To frame the sorts of formulas with positive occurrences of equality (modulo their prenex conjunctive normal form), we need some definitions which are similarly non-intuitive.
@@
We may loosen our terminology and call an arbitrary formula non-collapsible if its prenex conjunctive normal form is non-collapsible. We call a formula as such because they do not force the collapse of different levels of the hierarchy, that is, they are formulas whose truth do not require the identification of elements of different levels of the hierarchy. A more detailed explanation of the above definitions are in order. To offer it, we shall consider the context of a cumulative power of first level, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})$. As one may easily see, a literal which is not a positive equality is preserved in both directions between direct powers and their relative cumulative powers. Thus, the satisfaction of condition (i) of Definition 2.17 by a formula entails its preservation. Similarly, the satisfaction of condition (iii) of Definition 2.17 ensures that, in each conjunct of a formula in prenex conjunctive normal form, there is a subformula not containing positive equalities which is satisfied, so that we also obtain preservation. On the other hand, if a sentence satisfies condition (ii) Definition 2.17, there are three possible cases. If it satisfies condition (iii) of Definition 2.16, the presence of an existential quantifier on each side of the equality allows us to choose, for whatever choice of elements of $\mathcal{F}_{\beta}(\mathcal{A})$ on the other side, a suitable element of the same level as those, so that identity may be reached. A constant, on the other hand, does a similar job, since it forces the term in which occurs to be of the largest possible rank inside the cumulative power structure in question. If it satisfies condition (ii) of Definition 2.16, since if there is a universally bound variable on one side, then it also occurs on the other side, the ranks of the interpreted terms will always match, for they will be forced. it is possible to choose suitable elements of $\mathcal{F}_{\beta}(\mathcal{A})$ given any elements picked by the universal quantifiers. The constants, in turn, play a similar role, since they are of the highest rank possible in the relevant model, and similarly their values are of the highest possible rank which can be assumed. Thus, if constants are on one side of a positive equality and there are variables bound by existential quantifiers on the other side, we may match the ranks as required.
@@
We may see that cumulative powers do not, in general, preserve formulas containing positive occurrences of equality such that the existential quantifiers occur before the universal quantifiers bounding its variables. In fact, if universal quantifiers bound all of the variables of one side of equality, then such a sentence is not in general satisfied, because the levels of the chosen elements may not match. Consider for instance $\exists x_{1}...x_{n}\forall y_{1}...y_{m}\big(t(x_{1},...,x_{n})=s(y_{1},...,y_{m})\big)$, a formula which is not non-collapsible – that is, collapsible. Since $\rho\big(t^{\mathcal{F}_{1}(\mathcal{A})}(x_{1},...,x_{n})\big)=\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}$ and $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(y_{1},...,y_{m})\big)=\mathrm{max}\{\rho(y_{i})\}_{1\leq i\leq m}$, and since the universal quantifiers may pick elements in either $A$ or $\mathcal{F}^{*}_{1}(A)$, for a choice $b_{1},...,b_{m}\in A$ we would have $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=0$, which would require $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=0$. However, if $b_{1}\in\mathcal{F}^{*}_{1}(A)$, then $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=1$, which would mean $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=1$. Thus, there are no elements $a_{1},...,a_{n}\in\mathcal{F}_{1}(A)$ satisfying $\forall y_{1}...y_{m}\big(t(a_{1},...,a_{n})=s(y_{1},...,y_{m})\big)$ in $\mathcal{F}_{1}(\mathcal{A})$.
@@
Let $\mathcal{F}_{1}^{+}(\mathcal{A}$ be generated by the index set $I$, $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex disjunctive normal form. Then,
@@
We show the contrapositive by an induction on the number of quantifiers. Let $\varphi[a_{0},...,a_{n}]=\bigvee_{i\leq k}\bigwedge_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$ contain no quantifiers. If $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\varphi[a_{0},...,a_{n}]$, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\neg\psi_{i,l}[a_{0},...,a_{n}]$, then for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. If $\psi^{\prime}_{l,q}[a_{0},...,a_{n}]$ is a negated atom, we might observe, by our definitions, that means
@@
On the other hand, if $\psi^{\prime}_{l,q}[a_{0},...,a_{n}]$ is an atom, we may see, by our definitions, that means
@@
$\exists j\in I\big(\mathcal{A}\not\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i^{\prime},l}[\vartheta^{\delta}_{j}(a_{0}),...,\vartheta^{\delta}_{j}(a_{n})]\big)$.
@@
For the following results, we need the following notion. For each index set $I$ and $\sigma$-structure $\mathcal{A}$, we define the mapping $\vartheta^{-1}:\mathcal{F}_{1}(A)\to A^{I}$ such that, for each $a\in\mathcal{F}_{1}(A)$,
Lemma 2.20.
Proof.
@@
Suppose $\varphi$ satisfies condition (i). We proceed with an induction on the complexity of $\varphi$. If $\varphi$ is atomic, then it cannot be an equality, so the case is covered in both directions by Lemma 2.20. For the remaining cases, the non-trivial one is the existential. If $\mathcal{A}^{I}\models\forall x\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then for any $b\in A^{I}$, $\mathcal{F}_{1}^{+}(\mathcal{A})\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),b]$. Since $\vartheta^{-1}$ is surjective, that means for any $c\in\mathcal{F}_{1}(A)$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),\vartheta^{-1}(c)]$. Thus, by Lemma 2.18, $\varphi(x_{1},...,x_{n},y)$ is non-collapsible, so by the inductive hypothesis, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n},c]$. By the arbitrariness of $c$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\varphi[a_{0},...,a_{n}]$.
Suppose $\varphi$ satisfies condition (ii) because it satisfies condition (i) of Definition 2.16. We proceed with an induction on the number of quantifiers. Let $\mathcal{A}^{I}\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$ (which is possible if there are constants on each side of the positive equalities). That means for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{F}^{+}_{1}(\mathcal{A})\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$. If $\psi_{i^{\prime},l^{\prime}}$ is not a positive equality, then by Lemma 2.20, $\mathcal{A}^{I}\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. Otherwise, by assumption it must be of the form $t(x_{0},...,x_{n})=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$ for some constants $\mathrm{c}_{0},...,\mathrm{c}_{v}\in\sigma$. Then, $\mathcal{A}^{I}\models t[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. One may easily check, by our definitions, that implies $\mathcal{F}_{1}^{+}(\mathcal{A})\models t[a_{0},...,a_{n}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. By the arbitrariness of $i^{\prime}$ and $l^{\prime}$, we therefore have $\mathcal{F}_{1}^{+}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$. The inductive step is straightforward by Lemma 2.18 (1) and (2). If $\varphi$ satisfies condition (ii) of Definition 2.16, as above, the result may be shown by an induction on the number of quantifiers. The base case may be tackled by the same reasoning presented above for when $\varphi$ satisfies condition (i) (that is, going down to the level of literals and applying Lemma 2.20), and the argument for the inductive steps are the same. If now $\varphi$ satisfies condition (iii) of Definition 2.16, we may also deal with the base case and inductive steps in the same way.
@@
For $\beta\in\mathbf{On}$, let $\mathcal{F}_{\beta}(\mathcal{A})$ be generated by the index set $I$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex conjunctive normal form. If $\varphi$ is constant-free and non-collapsible, then for any $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$,
@@
Let $\mathcal{A}^{I}\models\varphi$. If $\varphi$ is a $\Sigma_{n}$ sentence $\exists x\Delta\psi(x)$, then there is $a\in A^{I}$ such that $\mathcal{A}^{I}\models\Delta\psi[a]$. Since $\vartheta^{-1}$ is surjective, there is $a^{\prime}\in\mathcal{F}_{1}(A)$ such that $a=\vartheta^{-1}(a^{\prime})$. Thus, by Lemma 2.21, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a^{\prime}]$, and so $\mathcal{F}^{+}_{1}(\mathcal{A})\models\exists x\Delta\psi(x)$. If $\psi$ is a $\Pi_{n}$ sentence $\forall x\Delta\psi(x)$, let $a\in\mathcal{F}_{1}(A)$. If $a\in A^{I}$, then $\mathcal{A}^{I}\models\Delta\psi[a]$, so by the same reasoning as before, by Lemma 2.21 we get $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a]$, since, in that case, $a=\vartheta^{-1}(a)$. If $a\not\in A^{I}$, that is, if $a\in A$, notice, by Lemma 2.18 (2), $\Delta\psi(x)$ is also non-collapsible. Thus, by Lemma 2.21, $\mathcal{F}_{1}(\mathcal{A})\models\Delta\psi[a]$ iff $\mathcal{A}^{I}\models\Delta\psi[\vartheta^{-1}(a)]$, which we know is the case. Therefore, by the arbitrariness of $a$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\Delta\psi(x)$. ∎

### 18 Mar 2026 v2

@@ 2.1. A concrete example
For $f\in\mathcal{F}^{*}_{1}(\mathbb{N})=\mathbb{N}^{\mathbb{N}}$, let $f^{+}$ be defined as $\forall j\in\mathbb{N}\big(f^{+}(j)=\mathrm{s}(f(j))\big)$. Consider the function $\mathrm{s}_{1}$, the extension of the successor function $\mathrm{s}$ to $\mathcal{F}_{1}(\mathbb{N})$:
@@
$\forall j\in\mathcal{F}_{1}(\mathbb{N})\big(a(j)=\mathrm{s}_{1}(j)+_{1}3\big)$,
@@
$a(j)=\begin{cases}j^{+}+_{1}3,\ \text{if}\ j\in\mathbb{N}^{\mathbb{N}}\\ \mathrm{s}(j)+3,\ \text{if}\ j\in\mathbb{N}\end{cases}$
@@
In order to have a visual representation of this sum we may take an order of $\mathcal{F}_{1}(\mathbb{N})$ such that $\mathbb{N}$ is the initial segment, and hence we may represent an element of $\mathcal{F}_{2}(\mathbb{N})$ by a column vector with the $j$-th coordinate being its value at $j$. Then the first vector below represents $\mathrm{s}_{1}$, and the sum $\mathrm{s}_{1}+_{2}3$ can be depicted as follows.
@@ 2.2. Preservation and properties of cumulative powers
In [12], Horn showed that any Horn sentence is preserved by direct products. In [4], Chang and Morel showed Horn sentences are also preserved by reduced products produced by the Fréchet filter, and as expressed in [7], Chang further generalized the result, showing that, in fact, any proper (that is, defined by proper filters) reduced product preserves Horn sentences.
A class of structures defined by a set of sentences is an elementary class, and if all those sentences are Horn sentences, call it a Horn class. The results mentioned above therefore say that any Horn class is closed by proper reduced (and thus direct) products. Chang and Morel showed in [4] that not every elementary class closed under direct products is a Horn class, but conjectured that any elementary class closed under reduced products is indeed a Horn class. In [13], Keisler proved the conjecture under the assumption the continuum hypothesis holds for at least one cardinal, whereas in [9], Galvin strengthened the result by eliminating that assumption. Briefly:
@@
Let the occurrence of an atom in a formula be positive if in the prenex conjunctive normal form of that formula the atom is preceded by an even number of negations, and let it be negative otherwise. As such, one might suspect the fragment preserved by cumulative powers is the class of Horn sentences without positive equalities, for, as we have seen, there is an imbalance in how the interpretation of constants, functions and other relations are defined in cumulative powers (that is, hereditarily), and how equality is defined (that is, as true identity, rather than hereditary identity). In that case, cumulative powers would preserve strictly fewer first-order sentences than reduced products. However, consider the sentence $\forall x\forall y\exists z(x+y=z)$. As one may check on our previous example of a hierarchy given by the natural numbers, that sentence is preserved and has a positive occurrence of equality. Therefore, the matter is not as simple. Furthermore, as we may comparatively see on the definitions of cumulative powers and direct powers, the former are much closer to the latter than to reduced products or reduced powers. Truth of formulas for functional elements of cumulative powers accord with how their truth is defined in direct powers. Therefore, if we want to characterize the fragment of first-order theory preserved by cumulative powers, our route should be to look at what fragment is preserved by direct powers.
In [20], Weinstein showed it is possible to give a – although not simple – characterization of that fragment not in terms of an inductively defined class of formulas, but as formulas arising from a given recursive decision procedure. To expose it, some definitions are needed.
@@
Let $\Phi$ be a finite autonomous set of formulas closed under $\bigwedge$ and $\bigvee$ up to equivalence. [^10] The definition for autonomous sets of formulas if offered in [3], pp. 425–426. It is, however, not of vital importance for the following results, so we omit it. Let $x$ be any variable and let $\varphi$, $\varphi^{\prime}$, and $\psi$ be formulas in $\bigvee\bigwedge\exists xF$. Fix $Q\in\{\exists,\forall\}$. Suppose for every disjunct $\varphi^{\prime}$ of $\varphi$ and every disjunct $\psi^{\prime}$ of $\psi$, there is a disjunct $\gamma^{\prime}$ of $\gamma$ such that for every conjunct $Qx\gamma^{\prime\prime}$ of $\gamma^{\prime}$, there are conjuncts $Qx\varphi^{\prime\prime}$ of $\varphi^{\prime}$ and $Qx\psi^{\prime\prime}$ of $\psi^{\prime}$ such that $\varphi^{\prime\prime}\times\psi^{\prime\prime}\Rightarrow\gamma^{\prime\prime}$. Then, $\varphi\times\psi\Rightarrow\gamma$. Moreover, if $\varphi$ and $\psi$ are elements of $\bigvee\bigwedge^{*}Qx\Phi$, then the supposition is necessary and sufficient for $\varphi\times\psi\Rightarrow\gamma$.
@@
Notice that $R$ is a computable predicate, as, given formulas $\varphi$, $\psi$ and $\gamma$, we can take $\Phi$ to be the set of atomic formulas occurring in $\varphi$, $\psi$ and $\gamma$, and it suffices to check for $n$ that is as large as the greatest among the quantifier ranks of those formulas.
@@
Notice that we may replace products with powers in the above corollary, that is:
@@
For the upcoming results, we need the following notion. For each index set $I$ and $\sigma$-structure $\mathcal{A}$, we define the mapping $\vartheta^{-1}:\mathcal{F}_{1}(A)\to A^{I}$ such that, for each $a\in\mathcal{F}_{1}(A)$,
Consider the following result.
Lemma 2.16.
Proof.
However, unlike other literals – and as we shall shortly see –, positive equalities are in general not preserved by cumulative powers. To frame those that are (modulo their prenex conjunctive normal form), we need some definitions.
@@
We may loosen our terminology and call an arbitrary formula non-collapsible if its prenex conjunctive normal form is non-collapsible. We call a formula as such because they do not force the collapse of different levels of the hierarchy, that is, they are formulas whose truth do not require the identification of elements of different levels of the hierarchy. A more detailed explanation of the above definitions is in order. To offer it, we shall consider the context of a cumulative power of first level, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})$. As one may easily see, a literal which is not a positive equality is preserved in both directions between direct powers and their relative cumulative powers. Thus, the satisfaction of condition (i) of Definition 2.18 by a formula entails its preservation. Similarly, the satisfaction of condition (iii) of Definition 2.18 ensures that, in each conjunct of a formula in prenex conjunctive normal form, there is a subformula not containing positive equalities which is satisfied, so that we also obtain preservation. On the other hand, if a sentence satisfies condition (ii) of Definition 2.18, there are three possible cases. If it satisfies condition (iii) of Definition 2.17, the presence of an existential quantifier on each side of the equality allows us to choose, for whatever choice of elements of $\mathcal{F}_{\beta}(\mathcal{A})$ on the other side, a suitable element of the same level as those, so that identity may be reached. A constant, on the other hand, does a similar job, since it forces the term in which it occurs to be of the largest possible rank inside the cumulative power structure in question. If it satisfies condition (ii) of Definition 2.17, since if there is a universally bound variable on one side, then it also occurs on the other side, the ranks of the interpreted terms will always match, for it is possible to choose suitable elements of $\mathcal{F}_{\beta}(\mathcal{A})$ given any elements picked by the universal quantifiers. The constants, in turn, play a similar role, since they are of the highest rank possible in the relevant model, and similarly their values are of the highest possible rank which can be assumed. Thus, if constants are on one side of a positive equality and there are variables bound by existential quantifiers on the other side, we may match the ranks as required.
@@
We may see that cumulative powers do not, in general, preserve formulas containing positive occurrences of equality such that the existential quantifiers occur before the universal quantifiers bounding its variables. In fact, if universal quantifiers bound all of the variables of one side of the equality, then such a sentence is not in general satisfied, because the levels of the chosen elements may not match. Consider for instance $\exists x_{1}...x_{n}\forall y_{1}...y_{m}\big(t(x_{1},...,x_{n})=s(y_{1},...,y_{m})\big)$, a formula which is collapsible. Since $\rho\big(t^{\mathcal{F}_{1}(\mathcal{A})}(x_{1},...,x_{n})\big)=\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}$ and $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(y_{1},...,y_{m})\big)=\mathrm{max}\{\rho(y_{i})\}_{1\leq i\leq m}$, and since the universal quantifiers may pick elements in either $A$ or $\mathcal{F}^{*}_{1}(A)$, for a choice $b_{1},...,b_{m}\in A$ we would have $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=0$, which would require $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=0$. However, if $b_{1}\in\mathcal{F}^{*}_{1}(A)$, then $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=1$, which would mean $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=1$. Thus, there are no elements $a_{1},...,a_{n}\in\mathcal{F}_{1}(A)$ satisfying $\forall y_{1}...y_{m}\big(t(a_{1},...,a_{n})=s(y_{1},...,y_{m})\big)$ in $\mathcal{F}_{1}(\mathcal{A})$.
@@
Let $\mathcal{F}_{1}^{+}(\mathcal{A})$ be generated by the index set $I$, $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex disjunctive normal form. Then,
@@
We show the contrapositive by an induction on the number of quantifiers. Let $\varphi[a_{0},...,a_{n}]=\bigvee_{i\leq k}\bigwedge_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$ contain no quantifiers. If $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\varphi[a_{0},...,a_{n}]$, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\neg\psi_{i,l}[a_{0},...,a_{n}]$, then for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. If $\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$ is a negated atom, we might observe, by our definitions, that means
@@
On the other hand, if $\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$ is an atom, we may see, by our definitions, that means
@@
$\exists j\in I\big(\mathcal{A}\not\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[\vartheta^{\delta}_{j}(a_{0}),...,\vartheta^{\delta}_{j}(a_{n})]\big)$.
@@
Suppose $\varphi$ satisfies condition (i). We proceed with an induction on the complexity of $\varphi$. If $\varphi$ is atomic, then it cannot be an equality, so the case is covered in both directions by Lemma 2.16. For the remaining cases, the non-trivial one is the existential. If $\mathcal{A}^{I}\models\forall x\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then for any $b\in A^{I}$, $\mathcal{A}^{I}\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),b]$. Since $\vartheta^{-1}$ is surjective, that means for any $c\in\mathcal{F}_{1}(A)$, $\mathcal{A}^{I}\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),\vartheta^{-1}(c)]$. Thus, by Lemma 2.19, $\varphi(x_{1},...,x_{n},y)$ is non-collapsible, so by the inductive hypothesis, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n},c]$. By the arbitrariness of $c$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\varphi[a_{0},...,a_{n}]$.
Suppose $\varphi$ satisfies condition (ii) because it satisfies condition (i) of Definition 2.17. We proceed with an induction on the number of quantifiers. Let $\mathcal{A}^{I}\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$ (which is possible if there are constants on each side of the positive equalities). That means for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{A}^{I}\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$. If $\psi_{i^{\prime},l^{\prime}}$ is not a positive equality, then by Lemma 2.16, $\mathcal{A}^{I}\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. Otherwise, by assumption it must be of the form $t(x_{0},...,x_{n})=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$ for some constants $\mathrm{c}_{0},...,\mathrm{c}_{v}\in\sigma$. Then, $\mathcal{A}^{I}\models t[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. One may easily check, by our definitions, that implies $\mathcal{F}_{1}^{+}(\mathcal{A})\models t[a_{0},...,a_{n}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. By the arbitrariness of $i^{\prime}$ and $l^{\prime}$, we therefore have $\mathcal{F}_{1}^{+}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$. The inductive step is straightforward by Lemma 2.19 (1) and (2). If $\varphi$ satisfies condition (ii) of Definition 2.17, as above, the result may be shown by an induction on the number of quantifiers. The base case may be tackled by the same reasoning presented above for when $\varphi$ satisfies condition (i) (that is, going down to the level of literals and applying Lemma 2.16), and the argument for the inductive steps are the same. If now $\varphi$ satisfies condition (iii) of Definition 2.17, we may also deal with the base case and inductive steps in the same way.
@@
For $\beta\in\mathbf{On}$, let $\mathcal{F}_{\beta}(\mathcal{A})$ be generated by the index set $I$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex conjunctive normal form. If $\varphi$ is constant-free and non-collapsible, then for any $a_{0},...,a_{n}\in\mathcal{F}_{\beta}(A)$,
@@
Let $\mathcal{A}^{I}\models\varphi$. If $\varphi$ is a $\Sigma_{n}$ sentence $\exists x\Delta\psi(x)$, then there is $a\in A^{I}$ such that $\mathcal{A}^{I}\models\Delta\psi[a]$. Since $\vartheta^{-1}$ is surjective, there is $a^{\prime}\in\mathcal{F}_{1}(A)$ such that $a=\vartheta^{-1}(a^{\prime})$. Thus, by Lemma 2.21, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a^{\prime}]$, and so $\mathcal{F}^{+}_{1}(\mathcal{A})\models\exists x\Delta\psi(x)$. If $\varphi$ is a $\Pi_{n}$ sentence $\forall x\Delta\psi(x)$, let $a\in\mathcal{F}_{1}(A)$. If $a\in A^{I}$, then $\mathcal{A}^{I}\models\Delta\psi[a]$, so by the same reasoning as before, by Lemma 2.21 we get $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a]$, since, in that case, $a=\vartheta^{-1}(a)$. If $a\not\in A^{I}$, that is, if $a\in A$, notice, by Lemma 2.19 (2), $\Delta\psi(x)$ is also non-collapsible. Thus, by Lemma 2.21, $\mathcal{F}_{1}(\mathcal{A})\models\Delta\psi[a]$ iff $\mathcal{A}^{I}\models\Delta\psi[\vartheta^{-1}(a)]$, which we know is the case. Therefore, by the arbitrariness of $a$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\Delta\psi(x)$. ∎